OpenStudy (some.random.cool.kid):

Which of the following is a quadratic equation that has the roots x = 2 and x = −7? x2 − 5x − 14 = 0 x2 − 5x = −14 x2 + 5x = 14 x2 + 5x = −14

OpenStudy (some.random.cool.kid):

I believe that this is C but would like some clarification. and yeah lunch seems great right now lol @Majesty69

OpenStudy (some.random.cool.kid):

@Maddy1251

OpenStudy (some.random.cool.kid):

i accidentally pressed the wrong name lol

OpenStudy (maddy1251):

Lol it happens! Do you know how to find roots?

OpenStudy (some.random.cool.kid):

mmm based off of context clues I could figure it out but not off the top of my head, no.

OpenStudy (maddy1251):

Roots are what makes a function equal 0, when we plug in certain 'x' terms, we get the equation to equal zero. Factoring is a great way to help, however in this case if you do not want to factor, you can plug in your given roots to find what yields 0.

OpenStudy (some.random.cool.kid):

oh also i dont believe its c i think B my bad. I just realized theers a negative but you can still explain. I want to know for sure.

OpenStudy (maddy1251):

I like all my equations to be set to zero when finding roots.

OpenStudy (maddy1251):

x2 − 5x − 14 = 0 Good. x2 − 5x = −14 --> x^2-5x+14=0 x2 + 5x = 14 --> x^2+5x-14=0 x2 + 5x = −14 ----> x^2+5x+14=0.

OpenStudy (maddy1251):

So, now we can just plug in x=2 and x=-7.

OpenStudy (some.random.cool.kid):

wow really ... I was considering that but im not sure if the zero plays a role, and if i does play a role i dont know how it does.

OpenStudy (maddy1251):

I will explain why we have zeros first, then proceed.

OpenStudy (maddy1251):

When we factor an equation like (x+7)(x-5)=0, we want to know how to get (0)(0)=0 In this case x=-7 for the first part, and x=5 for the second. These are known as the roots.

OpenStudy (maddy1251):

That's just an example, not sure if those roots truly work, but its a vision to see what roots really are and how they work.

OpenStudy (maddy1251):

Then when you would expand the problem, both of those roots would produce the polynomial to equal zero.

OpenStudy (maddy1251):

I have to eat real quick, but the answer is C. Try figuring it out to see why it is C.

OpenStudy (some.random.cool.kid):

a hah so far i understand. and ok will do.

OpenStudy (maddy1251):

I am back.

OpenStudy (some.random.cool.kid):

wow So i was right in the first place when I said C but then I second guess myself because of the negative. However I dont wish to guess that, think you can perhaps explain what happened to the negative.

OpenStudy (some.random.cool.kid):

does it have soemthing to do with cancelling out the numbers?

OpenStudy (some.random.cool.kid):

when I distribute?

OpenStudy (maddy1251):

No, and well you don't have to distribute. Not yet.

OpenStudy (maddy1251):

Remember how I set them to zero?

OpenStudy (some.random.cool.kid):

yeah

OpenStudy (maddy1251):

We have roots x=2 and x=-7

OpenStudy (maddy1251):

(x+2)(x-7)= ? Distribute that

OpenStudy (some.random.cool.kid):

ok give me a minute

OpenStudy (maddy1251):

x2 − 5x − 14 = 0 Good. x2 − 5x = −14 --> x^2-5x+14=0 x2 + 5x = 14 --> x^2+5x-14=0 x2 + 5x = −14 ----> x^2+5x+14=0. See which of these equals (x+2)(x-7) Take your time

OpenStudy (some.random.cool.kid):

the two x's make x^2 + x7 - 2x + 14? or was it a minus 14?

OpenStudy (some.random.cool.kid):

0.o a 5? where did that come from?

OpenStudy (maddy1251):

I will distribute for you.

OpenStudy (maddy1251):